# Vosh: Voxel-Mesh Hybrid Representation for Real-Time View Synthesis

### v1

@@ 4.2. Voxels to mesh conversion
Based on the above high-resolution voxel grid, we employ Marching Cubes (Lorensen and Cline, 1998) to extract an explicit mesh. However, surfaces derived solely from density information in volume rendering often exhibit roughness and are unsuitable for direct use in hybrid representation. Inspired by NeRF2Mesh (Tang et al., 2023), we adopt a synergistic approach involving differentiable rendering and surface operations to refine the explicit mesh.
The objective is to convert a subset of voxels that demonstrate similar rendering quality in both voxel and explicit mesh representations into a refined mesh. Typically, these converted voxels correspond to regions in the scene characterized by uncomplicated geometry and texture. The surface training specifically entails the coarse selection of voxels and subsequent surface refinement using differentiable rendering techniques.
Coarse selection of voxels. By analyzing the rendering results of NeRF2Mesh (Tang et al., 2023) (Figure 5) and BakedSDF (Yariv et al., 2023), we found that using mesh representation to reconstruct the unbounded background region far from the camera presents a formidable challenge. Achieving rendering quality comparable to voxels in this remote area is inherently demanding. Moreover, influenced by the contraction function, triangles extracted from voxels in the contraction region tend to enlarge and may break apart as they extend away from the camera. The presence of large triangles can adversely affect rendering quality, particularly on WebGL platforms, while introducing subdivision for these large triangles to enhance rendering quality incurs additional storage overhead.
Thus, we only convert voxels with surfaces located in the region near the camera into mesh. Criteria such as edge length and the number of triangles in isolated patches guide the decision on which explicit surfaces to retain. By default, we preserve the cube area with $90\%$ of the length of the bounding box in outdoor scenes. This conversion strategy, depicted in Figure 3, contrasts with indiscriminate conversion of voxels containing implicit surfaces into triangular meshes, which results in diminished rendering quality. So these explicit surfaces are further refined in the next differentiable appearance optimization.
Figure 3. The voxels to mesh conversion without (w/o) coarse selection of voxels obtains worse rendering results than with (w/) coarse selection, as shown by the diffuse of Vosh for voxel and mesh components respectively.
Differentiable surface refinement. We perform differentiable rasterization operations on the mesh to obtain intersections between rays and surfaces. Appearance information is obtained by querying the voxel representation based on the corresponding spatial positions. We also employ nvdiffrast (Laine et al., 2020) for differentiable rendering to optimize the appearance information of the mesh, while simultaneously tracking the rendering error corresponding to each face for surface optimization. It is worth noting that we do not enforce the mesh to represent the entire scene. Therefore, during differentiable rendering, we only calculate the loss for pixels corresponding to rays that intersect with the surface. We also perform face merging and mid-point subdivision operations based on the normal change of the faces and rendering errors. In contrast to NeRF2Mesh (Tang et al., 2023), which solely relies on rendering errors for operations, we additionally utilize the local normal change rate of faces as an optimization criterion. This enhances face density in regions with complex geometry and reduces storage overhead for faces in areas with simple geometry.
@@
Following the conversion from voxels to mesh, we have obtained an initial hybrid representation that includes both voxels and mesh. Nevertheless, recognizing significant potential for optimization in rendering quality and volume occupancy for both representations, we are committed to further refining the hybrid representation through further fine-tuning.
@@
Figure 4. The proposed hybrid rendering (c) integrates volume rendering (a) and surface rendering (b) for the Vosh.
4.3. Optimization
@@
After the mesh conversion and optimization stage, we obtain an explicit surface suitable for representing regions in the scene with simple geometry and textures, as well as a optimized high-resolution voxel grid. By combining these two components, we can obtain a coarse hybrid representation with a significant amount of redundancy between the components. This may result in low rendering efficiency and potential rendering conflicts. So we further perform hybrid rendering optimization and voxel adjustment based on the components for a more compact and efficient representation.
@@
(4) $\mathrm{\mathbf{C}}_{vosh}=\mathrm{\mathbf{C}}_{vosh,d}+\text{MLP}(\mathrm{\mathbf{C}}_{vosh,d},\mathrm{\mathbf{F}_{vosh},\mathrm{\mathbf{d}}}),$
(5) $\mathrm{\mathbf{F}}_{vosh}=\sum_{i}w_{i}\mathrm{\mathbf{f}}+w_{m}\mathrm{\mathbf{f}}_{m},$
(6) $\mathrm{\mathbf{C}}_{vosh,d}=\sum_{i}w_{i}\mathrm{\mathbf{c}}_{d,i}+w_{m}\mathrm{\mathbf{c}}_{d,m},$
@@
where $\mathrm{\mathbf{C}}_{vosh}$, $\mathrm{\mathbf{C}}_{vosh,d}$ and $\mathrm{\mathbf{F}}_{vosh}$ represent the color, diffuse color, and view-dependent features after hybrid rendering, respectively. $w_{m}=(1-\sum_{i}w_{i})$ represents the rendering weight for the mesh, which is the remaining portion of the cumulative voxel weights. $\mathrm{\mathbf{c}}_{d,m}$ and $\mathrm{\mathbf{f}}_{m}$ represent the diffuse color and view-dependent features at the intersection point where the ray hits the surface. Three different rendering methods are shown in Figure 4.
@@
Based on the optimized mesh, we attempt to dynamically adjust the voxel grid, forming a high-quality hybrid representation. Since the trained mesh already possesses the capability to represent regions with simple geometry and textures, we need to consider how to simplify the voxel representation for the entire scene. We focus on retaining voxels in regions with complex geometry and textures to better complement the rendering of the hybrid representation with the mesh. Inspired by PlenOctrees (Yu et al., 2021), we propose a voxel adjustment loss based on surface mesh weights. Specifically, we calculate the rendering weight of mesh $w_{m}$ for rays hitting the surface and force it close to 1. The implementation is as follows:
(7) $\mathcal{L}_{voxel}=\lambda_{voxel}\frac{1}{K}\sum_{k}(1-\text{exp}(w_{m}-1)),$
@@
where $\lambda_{voxel}$ represents the weight of the voxel adjustment loss, and $K$ is the number of samples taken in the rays hitting the surface.
@@
The number of retained voxels is an important factor in determining the balance between rendering quality and speed. Concretely, retaining more voxels can enhance rendering quality but slow down rendering speed, while retaining fewer voxels can accelerate rendering speed but compromise rendering quality. Setting different voxel adjustment weights to limit the volumetric density during ray marching is one of the methods to achieve different trade-offs between speed and quality. Additionally, voxels and mesh may overlap at the same spatial location, potentially affecting rendering quality and the volume of exported assets. To address this issue, we first calculate the occupancy grid of the meshes under the training poses. During voxel adjustment, we ignore the voxels in the mesh-occupied grid. We also adjust the ratio of sparse voxels in the hybrid representation by setting different resolutions $r_{mesh}$ for the mesh-occupancy grid.
Table 1. Quantitative evaluation of the proposed method and SOTA methods with 3 outdoor scenes and 4 indoor scenes from Mip-NeRF 360 dataset (Barron et al., 2022). The methods that can render in real-time on mobile phones are highlighted.
Outdoor Indoor GTX 1660Ti iPhone 14 Pro
Methods Representations PSNR ↑ SSIM ↑ LPIPS ↓ PSNR ↑ SSIM ↑ LPIPS ↓ FPS @ 1920*1080 FPS @ 344*707
NeRF MLP 21.46 0.458 0.515 26.84 0.790 0.370 - -
Mip-NeRF 360 MLP 25.92 0.747 0.244 31.72 0.917 0.179 - -
Instant-NGP Voxel 23.90 0.648 0.369 29.47 0.877 0.273 - -
SNeRG++${}_{1024}$ Voxel 24.11 0.625 0.377 27.32 0.802 0.276 9.0 -
MERF Voxel + Plane 24.41 0.678 0.298 27.80 0.855 0.271 9.7 25.3
BakedSDF Mesh 23.40 0.577 0.351 27.20 0.845 0.300 131.0 -
MobileNeRF Mesh 22.90 0.524 0.463 25.75 0.757 0.453 92.7 153.3
NeRF2Mesh Mesh 22.74 0.497 0.463 25.20 0.721 0.340 231.0 210.0
Vosh-Light Voxel + Mesh 23.63 0.586 0.392 27.02 0.792 0.294 79.7 70.0
Vosh-Base Voxel + Mesh 23.74 0.592 0.389 27.07 0.791 0.291 38.0 33.7
5. Real-Time Rendering
@@
The real-time rendering based on Vosh is realized as a JavaScript 3D web application, leveraging the capabilities of the MERF real-time renderer. The rendering process is also hybrid and structured into two passes: the rasterization pass and the ray marching pass. Both passes are employed to make rendering result consistency between training on the GPU and inference on the web.
@@
In the rasterization pass, meshes are uniformly shaded. The fragment shader takes charge of sampling appearance textures, incorporating diffuse color and view-dependent feature. The output of the rasterization pass includes corresponding diffuse, view-dependent features and depth map. These output is then used in the subsequent ray marching pass.
@@
In the ray marching pass, we also employ a multi-resolution hierarchical structure for the occupancy grid introduced in SNeRG (Hedman et al., 2021) to mark the position of occupied voxels. This structure creates binary mask images in multiple resolution levels through max-pooling. We use the multi-resolution occupancy grid to perform spatial skips during ray marching. Simultaneously, the depth map obtained in the rasterization pass further reduces the sampling distance.
@@
6. Experiments
We have implemented real-time view synthesis using the proposed hybrid representation Vosh, and carried out comprehensive evaluations encompassing both rendering quality and speed. We also compare our method with state-of-the-art (SOTA) methods, like offline classical methods (NeRF (Mildenhall et al., 2021), Mip-NeRF 360 (Barron et al., 2022), Instant-NGP (Müller et al., 2022)) and real-time methods (Mobile-Nerf (Chen et al., 2023), NeRF2Mesh (Tang et al., 2023), BakedSDF (Yariv et al., 2023), MERF (Reiser et al., 2023), SNeRG++ (Hedman et al., 2021)). As done by SOTA methods, we used the challenging unbounded dataset Mip-NeRF 360 (Barron et al., 2022) for evaluation. Next, we elaborate the details of the experimental results.
@@ 6. Experiments
Figure 5. The rendering results and zoomed-in images obtained by our method, as well as some SOTA methods based on voxels or mesh representation.
6.1. Rendering quality comparison
@@
We conduct both quantitative and qualitative evaluation of rendering quality for view synthesis obtained by the proposed method and SOTA methods. Quantitative evaluation uses the metrics of PSNR, SSIM (Wang et al., 2004) and LPIPS (Zhang et al., 2018) for the assessment of rendering quality. Table 1 shows the statistics of the quantitative evaluation of different methods.
Here, we classify these methods based on their ability to achieve real-time rendering on mobile devices like laptops and mobile phones. Methods such as SNeRG, MERF, and BakedSDF can only attain real-time rendering on select laptops with robust GPUs. It can be seen there exists an obvious gap between representations based on voxels and meshes. While BakedSDF narrows the quality gap with its millions of triangles, its substantial memory consumption makes it impractical for deployment on mobile phones.
@@
Vosh demonstrates superior rendering quality in real-time rendering on mobile devices, even surpassing some methods that cannot achieve real-time rendering on mobile devices. For the outdoor scenes, Vosh enhances rendering quality over mesh-based methods by leveraging sparse voxels through the use of a contraction function (see Figure 5, 6 and 7). However, for the indoor scenes, methods have limitations in the representational capacity of tiny view-dependent MLPs. While Vosh achieves real-time rendering on mobile devices by employing a tiny view-dependent MLP, it has slightly lower rendering quality in indoor scenes compared to MLP-based methods.
@@
We provide a qualitative comparison with SOTA methods based on voxels (Reiser et al., 2023) and mesh (Tang et al., 2023) on the examples in Figure 5. We select one distant view and one close-up view for local magnification for better exhibition. The mesh-based methods often lack the ability to represent background regions far from the camera location, whereas voxel-based methods necessitate higher computational and storage costs to achieve enhanced rendering quality. Contrarily, our method improves rendering speed while essentially maintaining the rendering quality comparable to that of voxel-based methods. We refer the reader to the companion video for visual demonstrations of the results.
Comparison with VMesh. It is worth noting that VMesh (Guo et al., 2023) is also built on the hybrid representation of volume and mesh, but primarily designed for rendering individual objects or scenes with boundaries, posing challenges in handling unbounded scenes like the examples in Figure 5 and 7. Nonetheless, our method produces better rendering results than VMesh, even for the synthetic bounded object as shown in Figure 6.
Figure 6. The rendering results obtained by ours and VMesh (Guo et al., 2023).
6.2. Rendering speed comparison

### v2

@@ IV-B Meshing density field
Differentiable surface refinement. Based on the above high-resolution voxel grid, we employ Marching Cubes [34] for the density field of initial voxel grid to extract an explicit mesh for evaluating the quality gap between voxel and mesh representation. However, surfaces meshed directly from volumetric density field often exhibit roughness that have to be improved for enhancing the rendering quality. Inspired by NeRF2Mesh [14], we adopt an ad-hoc scheme involving differentiable rendering and surface operations to refine the explicit mesh firstly.
Concretely, we perform differentiable rasterization operations on the mesh to obtain intersections between rays and surfaces. Appearance information is obtained by querying the voxel representation based on the corresponding spatial positions. Similar to NeRF2Mesh [14], we employ nvdiffrast [35] for differentiable rendering to optimize the appearance and the trainable offset of vertices, while simultaneously tracking the rendering error corresponding to each face for surface optimization.
In addition to rendering errors, we also calculate the normal angle between the adjacent faces of each face for performing mesh refinement and decimation, thereby improving the representation ability of the mesh and reducing memory usage. Specifically, for coarse mesh extracted based on initial voxel grid, we use differentiable rendering to learn offsets for each vertex of the mesh. Meanwhile, we calculate the rendering error $E_{render}$ and the normal angle between the adjacent faces $A_{normal}$ for each face. We set the rendering error threshold $e_{sub}$ and angle threshold $a_{sub}$ for face subdivision based on percentiles:
$e_{sub}=percentile(E_{render},90),$ (3)
$a_{sub}=percentile(A_{normal},95).$ (4)
Similarly, we set the rendering error threshold $e_{dec}$ and angle change threshold $a_{dec}$ for face simplification:
$e_{dec}=percentile(E_{render},50),$ (5)
$a_{dec}=percentile(A_{normal},10).$ (6)
For faces with $E_{render}>e_{sub}$ or $A_{normal}>a_{sub}$, we perform mid-point subdivision [36] to increase face density and enhance representation ability. For faces with $E_{render}<e_{dec}$ or $A_{normal}<a_{dec}$, we perform faces simplification [37] to improve rendering efficiency.
In contrast to NeRF2Mesh [14], which solely relies on rendering errors for operations, we additionally utilize the angle of local normal of faces as an optimization criterion. This enhances face density in regions with complex geometry and reduces storage overhead for faces in areas with simple geometry.
Figure 3: Visualization of conversion loss $E_{convert}$ based on rendering quality. (a) and (b) are the rendering images based on initial voxels and mesh with differentiable surface refinement, respectively. (c) is the gray-scale image of conversion loss $E_{convert}$ calculated based on the rendering quality of both representations. (d) is to filter out the top $10\%$ $E_{convert}$ in the space.
Meshing based on rendering quality. We can obtain the conversion error field $E_{convert}$ of different representations in the space based on differentiable rendering. Specifically, for voxel representation, we allocate rendering errors to corresponding positions based on the rendering weights of different sampling points. For mesh representation, the rendering errors at the corresponding position is recorded at the intersection of the surface and rays. We use grids to accumulate rendering errors for different representations and calculate the average rendering errors $E_{voxel}$ for voxel representation and $E_{mesh}$ for mesh representation. We calculate the L1 distance $E_{convert}=E_{mesh}-E_{voxel}$ of rendering errors in regions where neither is zero as the conversion errors, as shown in Figure 3.
It is observed in Figure 3 that the main distribution of $E_{convert}$ is in the background area far from the center of cameras. This is probably because 1) the reconstruction quality of the central area which is fully observed by cameras is higher enough in mesh representation; 2) contraction function leads to significant loss of information in background voxels. So the meshing operation is focused on the background area that lacks sufficient observation and is affected by the contraction function. For the scenes based on contraction function in $[-2,2]$, we define the central region as $[-0.5,0.5]$ and the outside as the background region. Then, we count voxels with background area and sort them, and also set the threshold in percentiles as $e_{remove}=percentile(E_{convert},p_{remove})$. When $p_{remove}$ is set to 0.9, the voxels with their $E_{convert}$ values greater than the threshold $e_{remove}$ is shown in Figure 3(d). Finally, we remove the faces converted from voxels located in the background area with $E_{convert}>e_{remove}$. The remaining mesh is the result of meshing the regions with small gap in rendering quality in the density field.
@@
Following the rendering quality based meshing process in density field, we have obtained an initial hybrid representation that includes both voxels and mesh. Nevertheless, considering significant potential for optimization in rendering quality and volume occupancy for both representations, we are committed to further refining the hybrid representation through the next optimization.
@@
Figure 4: The proposed hybrid rendering (c) integrates volume rendering (a) and surface rendering (b) for the Vosh.
TABLE I: Quantitative evaluation of the proposed method and SOTA methods on 3 outdoor scenes and 4 indoor scenes from Mip-NeRF 360 dataset [38]. The methods based on hybrid representation are listed separately below. means the method is not supported on the tested device.
Outdoor Indoor Intel Iris Xe 80EU iPhone 14 Pro
Methods PSNR ↑ SSIM ↑ LPIPS ↓ PSNR ↑ SSIM ↑ LPIPS ↓ FPS @ 800$\times$600 FPS @ 347$\times$600
NeRF 21.46 0.458 0.515 26.84 0.790 0.370 - -
Mip-NeRF 360 25.92 0.747 0.244 31.72 0.917 0.179 - -
Instant-NGP 23.90 0.648 0.369 29.47 0.877 0.273 - -
Zip-NeRF 26.87 0.807 0.176 32.03 0.925 0.167 - -
3DGS 26.40 0.805 0.173 30.41 0.920 0.190 - -
SNeRG++ 24.48 0.646 0.358 27.90 0.837 0.238 22.0 -
MERF 24.41 0.678 0.298 27.80 0.855 0.271 17.3 25.3
BakedSDF 23.40 0.577 0.351 27.20 0.845 0.300 21.0 -
MobileNeRF 22.90 0.524 0.463 25.75 0.757 0.453 59.3 153.3
NeRF2Mesh 22.74 0.497 0.463 25.20 0.721 0.340 138.7 210.0
Adaptive Shells 23.17 0.606 0.389 29.19 0.872 0.285 - -
VMesh - - - - - - - -
Vosh-Light 23.73 0.601 0.378 27.20 0.802 0.275 78.0 78.7
Vosh-Base 24.18 0.623 0.364 27.61 0.818 0.261 38.0 34.0
IV-C Optimization
@@
After meshing a portion of density field based on rending quality, we obtain an explicit surface suitable for representing regions in the scene with simple geometry and textures, as well as a optimized high-resolution voxel grid. By combining these two components, we can obtain a coarse hybrid representation with a significant amount of redundancy between the components. This may result in low rendering efficiency and potential rendering artifacts. So we further perform hybrid rendering optimization and voxel pruning based on the components for a more compact and efficient representation.
@@
$\mathrm{\mathbf{C}}_{vosh}=\mathrm{\mathbf{C}}_{vosh,d}+\text{MLP}(\mathrm{\mathbf{C}}_{vosh,d},\mathrm{\mathbf{F}_{vosh},\mathrm{\mathbf{d}}}),$ (7)
$\mathrm{\mathbf{F}}_{vosh}=\sum_{i}w_{i}\mathrm{\mathbf{f}}+w_{m}\mathrm{\mathbf{f}}_{m},$ (8)
$\mathrm{\mathbf{C}}_{vosh,d}=\sum_{i}w_{i}\mathrm{\mathbf{c}}_{d,i}+w_{m}\mathrm{\mathbf{c}}_{d,m},$ (9)
@@
where $\mathrm{\mathbf{C}}_{vosh}$, $\mathrm{\mathbf{C}}_{vosh,d}$ and $\mathrm{\mathbf{F}}_{vosh}$ represent the color, diffuse color, and view-dependent features after hybrid rendering, respectively. $w_{m}=(1-\sum_{i}w_{i})$ represents the rendering weight for the mesh, which is the remaining portion of the cumulative voxel weights. $\mathrm{\mathbf{c}}_{d,m}$ and $\mathrm{\mathbf{f}}_{m}$ represent the diffuse color and view-dependent features at the intersection point where the ray hits the surface. Three different rendering procedures are shown in Figure 4.
@@
Based on the optimized mesh, we attempt to dynamically prune the voxel grid, forming a high-quality hybrid representation. Since the trained mesh already possesses the capability to represent regions with simple geometry and textures, we need to consider how to simplify the voxel representation for the entire scene. We focus on retaining voxels in regions with complex geometry and textures to better complement the rendering of the hybrid representation with the mesh. Inspired by PlenOctrees [6], we propose a voxel pruning loss based on surface mesh weights. Specifically, we calculate the rendering weight of mesh $w_{m}$ for rays hitting the surface and force it close to 1. The implementation is as follows:
$\mathcal{L}_{voxel}=\lambda_{voxel}\frac{1}{K}\sum_{k}(1-\text{exp}(w_{m}-1)),$ (10)
@@
where $\lambda_{voxel}$ represents the weight of the voxel pruning loss, and $K$ is the number of samples taken before the rays hits the surface.
@@
The number of retained voxels is an important factor in determining the balance between rendering quality and speed. Concretely, retaining more voxels can enhance rendering quality but slow down rendering speed, while retaining fewer voxels can accelerate rendering speed but compromise rendering quality. Setting different voxel pruning weights to limit the volumetric density during ray marching is one of the methods to achieve different trade-offs between speed and quality. Additionally, voxels and mesh may overlap at the same spatial location, potentially affecting rendering quality and the volume of exported assets. To address this issue, we first calculate the occupancy grid of the meshes under the training poses and filter the voxels in the mesh-occupied grid out during rendering and baking. We also adjust the ratio of sparse voxels in the hybrid representation by setting different resolutions $r_{mesh}$ for the mesh-occupancy grid, which can also help suppress density of voxel component in Vosh.
V Real-Time Rendering
@@
The real-time rendering based on Vosh is realized as a JavaScript 3D web application, leveraging the capabilities of the MERF real-time renderer. The rendering process is also hybrid and structured into two passes: the rasterization pass and the ray marching pass. Both passes are employed to make the rendering result consistent between training on the GPU and inference on the web.
@@
In the rasterization pass, meshes are uniformly shaded. The fragment shader takes charge of sampling appearance textures, incorporating diffuse color and view-dependent feature. The output of the rasterization pass includes corresponding diffuse, view-dependent features and depth map. This output is then used in the subsequent ray marching pass.
@@
In the ray marching pass, we also employ a multi-resolution hierarchical structure for the occupancy grid introduced in SNeRG [8] to mark the position of occupied voxels. This structure creates binary mask images in multiple resolution levels through max-pooling. We use the multi-resolution occupancy grid to perform spatial skips during ray marching. Simultaneously, the depth map obtained in the rasterization pass further reduces the sampling distance.
@@
Figure 5: The rendering results and zoomed-in images in outdoor scenes obtained by our method, as well as some SOTA methods based on voxels or mesh representation.
VI Experiments
We have implemented real-time view synthesis using the proposed hybrid representation Vosh, and carried out comprehensive evaluations encompassing both rendering quality and speed. We also compare our method with state-of-the-art (SOTA) methods, like the classical offline methods (NeRF [1], Mip-NeRF 360 [38], Instant-NGP [11]) and real-time methods (MobileNerf [13], NeRF2Mesh [14], BakedSDF [16], MERF [18], SNeRG++ [8]). As done by SOTA methods, we use the challenging unbounded dataset Mip-NeRF 360 [38] for evaluation. Here, we specifically list the two most relevant hybrid representation based methods of Adaptive Shells [19] and VMesh [20] as shown in Table I. For Adaptive Shells, we obtain quantitative results from its paper, while VMesh mostly handles individual objects or bounded scenes, while our method focuses more on real scenes. Next, we elaborate the details of the experiments.
Figure 6: More rendering results and zoomed-in images in indoor scenes obtained by our method, as well as some SOTA methods based on voxels or mesh representation.
VI-A Implementation details
In the stage of training initial voxels, we use the classic MSE loss to measure pixel-level color differences. We use the same optimizer settings as MERF, e.g., 25000 steps for optimization and initially setting learning rates to 1e-2, which warmup on the first 100 steps and ultimately decay to 1e-3. We use 1024 as the resolution of the marching cubes and clean mesh with similar operations in NeRF2Mesh.
In the stage of meshing density field, we use 20000 steps for mesh refinement. For the process of meshing density field, we set $p_{remove}=0.9$ for Vosh-Base and $p_{remove}=1.0$ for Vosh-Light. After the meshing stage, some messy floating faces may be obtained. We use similar parameters to repeat the operation of cleaning the mesh in the previous stage. And we encourage voxels to fill possible holes in the mesh by using L2 regularization terms with rendering weights for background colors.
There are 20000 steps for Vosh optimization. For the Vosh-Base model, we set voxel pruning loss $\lambda_{voxel}=0.001$ and disable the binary mesh-occupancy grid. For the Vosh-Light model, we use voxel pruning loss $\lambda_{voxel}=0.01$ and the resolution $r_{mesh}=128$ for mesh occupancy grid as hyper-parameters.
Besides, we did all experiments on a single NVIDIA RTX 4090. Three training stages take approximately 66, 35 and 58 minutes respectively. During the training process, the peak video memory usage for optimizing trainable parameters is about 11 Gigabytes. It should be noted though the training phase requires high-performance hardware, the inference phase can be performed directly on mobile devices with limited resources.
VI-B Rendering quality comparison
@@
We conduct both quantitative and qualitative evaluation of rendering quality for view synthesis obtained by the proposed method and SOTA methods. Quantitative evaluation uses the metrics of PSNR, SSIM [39] and LPIPS [40] for the assessment of rendering quality. Table I shows the statistics of the quantitative evaluation of different methods.
@@
Vosh demonstrates superior rendering quality in real-time rendering on mobile devices, even surpassing some methods that cannot achieve real-time rendering on mobile devices. For the outdoor scenes, Vosh enhances rendering quality over mesh-based methods by leveraging sparse voxels through the use of a contraction function (see Figure 5 and 6). Rendering quality of Vosh in outdoor scenes also exceeds that of Adaptive Shells, which is also based on hybrid representation. The main reason is that Adaptive Shells constrain volume rendering to the regions near the surface, while in outdoor scenes, pure volume rendering is mostly of necessity. However, the limitations in representational capacity of the methods using tiny view-dependent MLPs limit rendering quality in indoor scenes. While Vosh achieves real-time rendering on mobile devices by employing a tiny view-dependent MLP, it has slightly lower rendering quality in indoor scenes compared to MLP-based methods.
@@
We provide a qualitative comparison with SOTA methods based on voxels [18] and mesh [14] on the examples in Figure 5. We select one distant view and one close-up view with local magnification for better exhibition. The mesh-based methods often lack the ability to represent background regions far from the camera location, whereas volume-based methods necessitate higher computational and storage costs to achieve enhanced rendering quality. Contrarily, our method improves rendering speed while essentially maintaining the rendering quality comparable to that of volume-based methods. We refer the reader to the companion video for dynamic demonstrations of the results.
VI-C Rendering speed comparison